# Learning to Use AI for Learning: How Can We Effectively Teach and Measure Prompting Literacy for K–12 Students?

### v1

@@ Abstract
As Artificial Intelligence (AI) becomes increasingly integrated into daily life, there is a growing need to equip the next generation with the ability to apply, interact with, evaluate, and collaborate with AI systems responsibly. Prior research highlights the urgent demand from K-12 educators to teach students the ethical and effective use of AI for learning. To address this need, we designed an Large-Language Model (LLM)-based module to teach prompting literacy. This includes scenario-based deliberate practice activities with direct interaction with intelligent LLM agents, aiming to foster secondary school students’ responsible engagement with AI chatbots. We conducted two iterations of classroom deployment in 11 authentic secondary education classrooms, and evaluated 1) AI-based auto-grader’s capability; 2) students’ prompting performance and confidence changes towards using AI for learning; and 3) the quality of learning and assessment materials. Results indicated that the AI-based auto-grader could grade student-written prompts with satisfactory quality. In addition, the instructional materials supported students in improving their prompting skills through practice and led to positive shifts in their perceptions of using AI for learning. Furthermore, data from Study 1 informed assessment revisions in Study 2. Analyses of item difficulty and discrimination in Study 2 showed that True/False and open-ended questions could measure prompting literacy more effectively than multiple-choice questions for our target learners. These promising outcomes highlight the potential for broader deployment and highlight the need for broader studies to assess learning effectiveness and assessment design.
Introduction
@@
The rising coverage of AI in everyday life techniques for human interaction with AI (Van Brummelen, Shen, and Patton 2019). As AI products have increased adaptability and ease of use, students can access them easily in different contexts, including at home and school (Yim and Su 2024). Some AI commercial products, like ChatGPT, have been introduced to K-12 educational settings, aiming to personalize student learning experience (Zhang and Tur 2024; Nayak et al. 2023). Apart from its potential values, there are growing concerns about utilizing generative AI chatbot technologies including increasing cheating (Lee et al. 2024; Li et al. 2024; Mah et al. 2024) and the low quality of information retrieval in learning settings (Wang et al. 2023).
@@
While AI chatbots hold promise for K-12 education, few instructional interventions, to our knowledge, focus on teaching prompting literacy to K-12 students. Prompts are natural language instruction inputs that are used to communicate with AI chatbot technologies (Lo 2023). Prompting literacy includes helping students understand what AI can do to support learning, when to use AI effectively, and how to craft prompts for different types of assistance. While the potential of LLMs is promising, their current effectiveness remains relies heavily on the quality of the prompts they receive (Ekin 2023). By designing and engineering them carefully, prompts can play a key role in shaping the AI’s responses and guiding the output of LLMs to provide desired results (Lo 2023; White et al. 2023). As prompts become an important end-user communication channel with AI chatbots, scholars in various educational domains argue that the ability to create effective prompts that generate desired information is now an essential skill for students (Denny, Kumar, and Giacaman 2023; Korzynski et al. 2023; Meskó 2023; Woo, Guo, and Susanto 2024). Therefore, equipping students with prompting literacy is crucial to help students effectively communicate with AI chatbots during learning.
@@
This work presents a web-based interactive instructional system to improve secondary-education (middle- and high-school) students’ prompting literacy. It applies active learning (Bonwell and Eison 1991) and experiential learning (Ng et al. 2023) methods to engage students with prompting practice in three hypothetical learning scenarios. After a student writes a prompt, an AI auto-grader evaluates key dimensions and delivers immediate, detailed feedback. We hypothesized that practicing prompt writing in this way could help students learn to create effective prompts, supporting learning and confidence in AI use. We first outlined the module design. Then, we reported our classroom evaluation results. Specific research questions were listed under each evaluation study. Finally, we proposed ways to improve prompting literacy instruction.
Related Work
Prompting literacy education in K-12 classrooms
@@
It is believed that the ability to understand and use AI (or the lack thereof) will fuel the next digital divide in education (Trucano 2023). The emergence of generative AI makes such technologies more accessible to the general public and underscores that it is more urgent than ever to regulate the responsible and effective use of such technologies (Williamson, Molnar, and Boninger 2024). While K-12 teachers expressed their timely needs for AI competency materials for their students, especially prompting instructions (Lozano and Blanco Fontao 2023), the majority of current K-12 AI Literacy implementations emphasize understanding AI technologies than responsible use of AI. Additionally, existing AI competency frameworks primarily target adults in the workforce (Meskó 2023; Tseng et al. 2024; Zamfirescu-Pereira et al. 2023) and higher-education (Denny, Kumar, and Giacaman 2023; Giray 2023). As a focused area within AI literacy, prompting literacy has gained increasing attention since 2023 (Hwang, Lee, and Shin 2023; Lo 2023; Maloy and Gattupalli 2024; Dennison et al. 2024; Xiao et al. 2025). However, the lack of emphasis on K-12 learners persists, as few studies specifically target secondary education students, despite growing concerns about student AI use in K–12 education (Li et al. 2024; Uanachain and Aouad 2025).
Learning Sciences Foundations: Learn-by-Doing with Elaborated Immediate Feedback
@@
This study’s instructional design is grounded in two foundational learning sciences principles: learning-by-doing (Schank, Berman, and Macpherson 2013) and elaborated immediate feedback (Wang et al. 2019). Learning-by-doing emphasizes active engagement through direct interaction with tasks. Prior research has shown that this approach is more effective than passive methods (e.g., reading, video-watching) in online learning environments (Koedinger et al. 2015). While learning-by-doing has been applied in prompting literacy education (e.g., Promptly (Denny et al. 2023)), such efforts often lack elaborated immediate feedback, which may hinder effective learning or even lead to drop-out (Vasilyeva, De Bra, and Pechenizkiy 2008).
@@
Recent advances in large language models (LLMs) offer new opportunities to address this challenge. LLMs have demonstrated strong performance in rubric-based grading across domains (Henkel et al. 2024; Nguyen et al. 2023) and are capable of generating personalized feedback for open-ended responses (Xiao, Hou, and Stamper 2024; Nguyen et al. 2023). For example, GPT-4 achieves up to 85% grading accuracy on short-answer questions using zero-shot prompting strategies (Henkel et al. 2024). Building on these capabilities, our work integrates an AI auto-grading pipeline with prompt-based practice to close the gap in prompting literacy instruction—enabling learning-by-doing accompanied by elaborated, immediate feedback to support more efficient and effective learning.
Prompting Literacy Module Overview
Learning Objectives
@@
Based on the real needs of teachers (Lee et al. 2024) for instructions to teach students how to use AI for learning and the AI4K12 guideline (Touretzky et al. 2019), three learning objectives were derived: AI Capacity for Supporting Learning, Contexts to Use AI for Learning, and Effective Prompt Formation (Figure 2).
Learning Materials
@@
The main practice is conducted using three hypothetical learning scenarios. Each practice has four steps (Figure 1): 1) introduce the detailed scenario to students, including the domain, the context, the student’s current knowledge level, the learning goal the student is trying to achieve, and the resources the student has; 2) the student creates a prompt and clicks ”generate” to receive answers from an AI chatbot. This mimics the authentic interaction with general AI chatbots; 3) the student receives an answer from the AI chatbot based on the written prompt; 4) they can then click ”Check my prompt” to receive an auto-graded result and a detailed explanation. This feedback is provided based on the preset evaluation dimensions for each scenario (Table 1). After receiving the feedback, students can back to craft a new prompt for the same practice or move on to the next practice scenario. Detailed practice pipeline is in Figure 1.
@@
The three scenarios were designed under three secondary school subjects: biology, geography, and math. Each domain was paired with a unique instructional activity, such as ”broaden knowledge beyond course requirements”, ”prepare for a quiz tomorrow”, and ”struggle with an assignment due tomorrow”. Here is an example scenario:
@@
Automatic Grading Dimensions
In our study, the requirements for a high-quality prompt vary depending on the instructional activity the student is pursuing (Table 1). For example, a good prompt in the homework struggle situation should not explicitly ask for a direct answer. However, this criterion does not apply to other scenarios like course content extension. Therefore, when designing the criteria for a high-quality prompt, each practice for each scenario has specific dimensions to meet. Also, each dimension has its own in-context descriptions rooted in the general definitions. In this study, we created our own grading dimensions based on the CLEAR framework for prompt creation (Lo 2023) and other existing prompting recommendations (Meskó 2023). Based on this grading rubric, for each student-written prompt, AI automatic grading needed to provide both a True/False categorical pass situation and a detailed explanation for this decision for each included dimension in that practice task. In addition, it was required to provide an detailed explanation for each dimension.
Table 1: Grading dimensions and associated tasks for student-written prompts. ”✓” means the LLM assesses if the student’s prompt meets the definition with a binary response (Yes or No) and then gives corresponding grading explanations.
@@
After the practice, the student re-answered a 5-point Likert-scale survey question about ”I know how to use AI to help me learn” and the six multiple-choice questions again as the post-test. At the end, the student answered 6 open-ended survey questions to reflect on the learning experiences. The LLM used was OpenAI GPT-4o.
Assessment Question Design
Scenario-based assessments have been applied to measuring skills in varied domains, such as writing (Zhang et al. 2019), mathematics (Cayton-Hodges et al. 2012) and science (Bergner and von Davier 2019). Such situated assessment experience can engage learners and facilitate deep thinking by presenting realistic, contextual problems that promote critical thinking and knowledge application (Clark 2010; Herrington and Oliver 2000). Furthermore, as our instructional goals were competency-based, scenario-based assessments are suitable to test students’ abilities. Multiple Choice Questions (MCQs) enable automatic grading while providing consistent evaluations across diverse learner groups. When carefully crafted, MCQs can target any cognitive level in Bloom’s Taxonomy (Amer 2006), making them a versatile and effective component of our assessment approach. Therefore, the assessment consisted of six MCQs. Each question presents three answer choices, with only one being correct. In Study 1, we chose MCQs as the only assessment question type and situated them into different scenarios based on the learning objectives.
Data Collection and Analysis
In Study 1, we collected both students’ written prompts and corresponding AI-grading results. To answer RQ1, we extracted students’ last written prompt and graded them manually to assess AI auto-grading. After receiving 483 unique student-written prompts, two researchers used the same rubric to grade the students’ written prompts (Table 1). They first randomly selected 15% of the prompt data and conducted human grading based on the rubric. After that, the two researchers met to address the conflicts and iteratively refined the detailed grading book. Then they randomly selected another 15% of the prompt data and both two researchers graded it again, the inter-rater reliability for this round reached higher than 0.92 (almost perfect agreement following (McHugh 2012)). Then the two researchers met again to resolve the conflicts and graded the rest. We used human grading results for RQ2. For RQ3, we conducted a thematic analysis on student answers. One hundred and thirty-one students finished answering the self-reflection questions. Two researchers conducted a coding book iteratively to analyze their answers and reported the main themes.
Table 2: Grading Accuracy of the Auto-Grader
@@
The auto-grader achieved a high accuracy (higher than 0.9) in Relevance, Background, and Elaboration dimensions. For the inaccurate cases in these dimensions, one pattern is that the auto-grader tended to weigh heavily on some keywords but ignored other keywords (Nguyen et al. 2023). For instance, one prompt was ”You are preparing for a geography exam and have just learned the basic concepts of the water cycle Answer in Sanskrit.” Auto-grader gave it a False in Relevance and explained as ”The core of the problem lies in language choice and has nothing to do with the specific content of the water cycle.”.
@@
The accuracy for Conciseness and No Direct Answer was 0.93 and 0.88. The conflicted cases for Conciseness mostly included two types. First, the auto-grader focused on spelling errors and therefore graded them as False, while human graders were more tolerant about this. Second, prompts containing background information might be seen as non-concise by the auto-grader. For example, one student-written prompt was ”Can you tell me in detail about your extracurricular knowledge about cells besides the definition of cells?”. The auto-grader graded Conciseness as False and explained it as ”The question is relatively long and the sentences are not concise enough. It could be more refined.”
@@
As for No Direct Answer, sometimes the student asked a general question about ”How to solve a two-variable linear equation?”, but the auto-grader treated this as the student tried to ask for a direct answer indirectly and labeled it as False. On the other hand, sometimes the student asked for ”list the steps to solve this equation 10x+4y=3,-2x+10y = 4”, but auto-grader thought this prompt ”did not directly ask about the values of x and y, but rather asks for a solution.” and gave it True in No Direct Answer instead.
@@
Auto-grader received the lowest accuracy in Purpose dimension (0.85). The main reason for false positive cases in this dimension is that the auto-grader might over-generate the keywords in a student-written prompt, and consider that as providing a clear purpose. For instance, it counted ”I am a junior high school student. Please help me solve this problem using junior high school knowledge.” as True in Purpose dimension and explained it as ”the student’s goal is to solve this problem.” However, simply asking for ”solving this problem” without specifying the type or domain does not align with our rubric (”specific and clear purpose, e.g. explain the math concepts involved”).
@@
For true negative cases in Purpose, sometimes the auto-grader mixed the criteria of Purpose and No Answer, and therefore considered those involving asking for an answer as False. Another reason is that auto-grader sometimes requires too detailed purpose, which is outside of our targeted scope. For example, one explanation to an AI-graded False case (”I want to know about DNA”) is ”Although DNA is mentioned in this question, the purpose is not clear. Students should be more specific about what they want to learn about DNA.” However, as the scenario is novice learners who only finished the first biology class about cells, the auto-grader should not expect the prompt purpose to be that detailed.
@@
Students answered four 5-scale Likert questions as a pre-survey before doing the pre-test. After finishing the learning activities, they answered the last question in the pre-survey again. Specific questions and the student response distribution are shown in Fig.3.
@@
(Applied human-graded scores, Q2 0.93 0.65 0.94 0.12 0.59 - -
Yes-1/No-0) Q3 0.89 0.75 0.95 0.11 0.52 0.64 -
@@
Ninety-eight students completed both the pre- and post-test. However, no significant increment (p = .377) was found in the post-test (Mean = 4.4, Median = 5.0, SD = 0.96, full score = 6.0) due to the ceiling effect in the pre-test (Mean = 4.4, Median = 5.0, SD = 1.04). In other words, students had the basic skills to identify the prompt quality from a conceptual standpoint, although their prompt writing skills remain vary.

### v2

# Learning to Use AI for Learning: Teaching Responsible Use of AI Chatbot to K-12 Students Through an AI Literacy Module

@@ Abstract
As Artificial Intelligence (AI) becomes increasingly integrated into daily life, there is a growing need to equip the next generation with the ability to apply, interact with, evaluate, and collaborate with AI systems responsibly. Prior research highlights the urgent demand from K-12 educators to teach students the ethical and effective use of AI for learning. To address this need, we designed a Large-Language Model (LLM)-based module to teach prompting literacy. This includes scenario-based deliberate practice activities with direct interaction with intelligent LLM agents, aiming to foster secondary school students’ responsible engagement with AI chatbots. We conducted two iterations of classroom deployment in 11 authentic secondary education classrooms, and evaluated 1) AI-based auto-grader’s capability; 2) students’ prompting performance and confidence changes towards using AI for learning; and 3) the quality of learning and assessment materials. Results indicated that the AI-based auto-grader could grade student-written prompts with satisfactory quality. In addition, the instructional materials supported students in improving their prompting skills through practice and led to positive shifts in their perceptions of using AI for learning. Furthermore, data from Study 1 informed assessment revisions in Study 2. Analyses of item difficulty and discrimination in Study 2 showed that True/False and open-ended questions could measure prompting literacy more effectively than multiple-choice questions for our target learners. These promising outcomes highlight the potential for broader deployment and highlight the need for broader studies to assess learning effectiveness and assessment design.
1 Introduction
@@
The increasing presence of AI in everyday life has amplified the need for effective techniques for human–AI interaction, including within educational settings (van2019popstar). As AI products have increased adaptability and ease of use, students can access them easily in different contexts (hou2025exploring), including at home and school (yim2024artificial). Some AI commercial products, like ChatGPT, have been introduced to K-12 educational settings, aiming to personalize student learning experience (zhang2024systematic; nayak2023teaching). Apart from its potential values, there are growing concerns about utilizing generative AI chatbot technologies including increasing cheating (lee2024cheating; li2025unseen; xiao2025teachers) and the low quality of information retrieval in learning settings (wang2023unleashing; kazemitabaar2023novices).
@@
While AI chatbots hold promise for K-12 education, few instructional interventions, to our knowledge, focus on teaching prompting literacy to K-12 students. Prompts are natural language instruction inputs that are used to communicate with AI chatbot technologies (lo2023clear). Prompting literacy includes helping students understand what AI can do to support learning, when to use AI effectively, and how to craft prompts for different types of assistance. While the potential of LLMs is promising, their current effectiveness remains heavily dependent on the quality of the prompts they receive (ekin2023prompt). By designing and engineering them carefully, prompts can play a key role in shaping the AI’s responses and guiding the output of LLMs to provide desired results (lo2023clear; white2023prompt). As prompts become an important end-user communication channel with AI chatbots, scholars in various educational domains argue that the ability to create effective prompts that generate desired information is now an essential skill for students (denny2023conversing; korzynski2023artificial; mesko2023prompt; woo2024exploring). Therefore, equipping students with prompting literacy is crucial to help students effectively communicate with AI chatbots during learning.
@@
This work presents a web-based interactive instructional system to improve secondary-education (middle- and high-school) students’ prompting literacy. It applies active learning (bonwell1991active) and experiential learning (ng2023ai) methods to engage students with prompting practice in three hypothetical learning scenarios. After a student writes a prompt, an AI auto-grader evaluates key dimensions and delivers immediate, detailed feedback. We hypothesized that practicing prompt writing in this way could help students learn to create effective prompts, supporting learning and confidence in AI use. We first outlined the module design. Then, we reported our classroom evaluation results. Specific research questions were listed under each evaluation study. Finally, we proposed ways to improve prompting literacy instruction.
2 Related Work
2.1 Prompting literacy in K-12 classrooms
@@
It is believed that the ability to understand and use AI (or the lack thereof) will fuel the next digital divide in education (trucano2023ai). The emergence of generative AI makes such technologies more accessible to the general public and underscores that it is more urgent than ever to regulate the responsible and effective use of such technologies (williamson2024time). While K-12 teachers expressed their timely needs for AI competency materials for their students, especially prompting instructions (lozano2023education), the majority of current K-12 AI Literacy implementations emphasize understanding AI technologies than responsible use of AI. Additionally, existing AI competency frameworks primarily target adults in the workforce (mesko2023prompt; tseng2024assessing; zamfirescu2023johnny) and higher-education (denny2023conversing; giray2023prompt). As a focused area within AI literacy, prompting literacy has gained increasing attention since 2023 (hwang2023prompt; lo2023clear; maloy2024prompt; dennison2024consumers; xiao2025improving). However, the lack of emphasis on K-12 learners persists, as few studies specifically target secondary education students, despite growing concerns about student AI use in K–12 education (li2025unseen; uanachain2025generative).
2.2 Learn-by-Doing with Elaborated Immediate Feedback
@@
This study’s instructional design is grounded in two foundational learning sciences principles: learning-by-doing (schank2013learning) and elaborated immediate feedback (wang2019elaborated). Learning-by-doing emphasizes active engagement through direct interaction with tasks. Prior research has shown that this approach is more effective than passive methods (e.g., reading, video-watching) in online learning environments (koedinger2015learning). While learning-by-doing has been applied in prompting literacy education (e.g., Promptly (denny2023promptly)), such efforts often lack elaborated immediate feedback, which may hinder effective learning or even lead to drop-out (vasilyeva2008immediate).
@@
Recent advances in large language models (LLMs) offer new opportunities to address this challenge. LLMs have demonstrated strong performance in rubric-based grading across domains (henkel2024can; nguyen2023evaluating) and are capable of generating personalized feedback for open-ended responses (xiao2024exploring; nguyen2023evaluating). For example, GPT-4 achieves up to 85% grading accuracy on short-answer questions using zero-shot prompting strategies (henkel2024can). Building on these capabilities, our work integrates an AI auto-grading pipeline with prompt-based practice to close the gap in prompting literacy instruction—enabling learning-by-doing accompanied by elaborated, immediate feedback to support more efficient and effective learning.
3 Prompting Literacy Module Overview
3.1 Learning Objectives
@@
Based on the real needs of teachers (lee2024cheating) for instructions to teach students how to use AI for learning and the AI4K12 guideline (touretzky2019special), three learning objectives were derived: AI Capacity for Supporting Learning, Contexts to Use AI for Learning, and Effective Prompt Formation (Figure 2).
3.2 Learning Materials
@@
The main practice is conducted using three hypothetical learning scenarios. Each practice has four steps (Figure 1): 1) Introduce the detailed scenario to students, including the domain, the context, the student’s current knowledge level, the learning goal the student is trying to achieve, and the resources the student has. 2) The student creates a prompt and clicks “generate” to receive answers from an AI chatbot. This mimics the authentic interaction with a general-purpose AI chatbot. 3) The student receives an answer from the AI chatbot based on the written prompt. 4) They can then click “Check my prompt” to receive an auto-graded result and a detailed explanation. This feedback is provided based on the preset evaluation dimensions for each scenario (Table 1). After receiving feedback, students can go back to craft a new prompt for the same practice or move on to the next practice. The detailed practice pipeline is shown in Figure 1.
@@
The three scenarios were designed under three secondary school subjects: biology, geography, and math. Each domain was paired with a unique instructional activity, such as “broaden knowledge beyond course requirements”, “prepare for a quiz tomorrow”, and “struggle with an assignment due tomorrow”. Here is an example scenario:
@@
3.3 Automatic Grading Dimensions
In our study, the requirements for a high-quality prompt vary depending on the instructional activity the student is pursuing (Table 1). For example, a good prompt in the homework struggle situation should not explicitly ask for a direct answer. However, this criterion does not apply to other scenarios like course content extension. Therefore, when designing the criteria for a high-quality prompt, each practice for each scenario has specific dimensions to meet. Also, each dimension has its own in-context descriptions rooted in the general definitions. In this study, we created our own grading dimensions based on the CLEAR framework for prompt creation (lo2023clear) and other existing prompting recommendations (mesko2023prompt). Based on this grading rubric, for each student-written prompt, AI automatic grading needed to provide both a True/False categorical pass situation and a detailed explanation for this decision for each included dimension in that practice task. In addition, it was required to provide an detailed explanation for each dimension.
@@ 3.3 Automatic Grading Dimensions
Table 1: Grading dimensions and associated tasks for student-written prompts. “✓” means the LLM assesses if the student’s prompt meets the definition with a binary response (Yes or No) and then gives corresponding grading explanations.
@@ 4.1 Study Procedure
After the practice, the student re-answered a 5-point Likert-scale survey question about “I know how to use AI to help me learn” and the six multiple-choice questions again as the post-test. At the end, the student answered 6 open-ended survey questions to reflect on the learning experiences. The LLM used was OpenAI GPT-4o.
4.2 Assessment Question Design
Scenario-based assessments have been applied to measuring skills in varied domains, such as writing (zhang2019scenario), mathematics (cayton2012technology) and science (bergner2019process). Such situated assessment experience can engage learners and facilitate deep thinking by presenting realistic, contextual problems that promote critical thinking and knowledge application (clark2010formative; herrington2000instructional). Furthermore, as our instructional goals were competency-based, scenario-based assessments are suitable to test students’ abilities. Multiple Choice Questions (MCQs) enable automatic grading while providing consistent evaluations across diverse learner groups. When carefully crafted, MCQs can target any cognitive level in Bloom’s Taxonomy (amer2006reflections), making them a versatile and effective component of our assessment approach. Therefore, the assessment consisted of six MCQs. Each question presents three answer choices, with only one being correct. In Study 1, we chose MCQs as the only assessment question type and situated them into different scenarios based on the learning objectives.
4.3 Data Collection and Analysis
In Study 1, we collected both students’ written prompts and corresponding AI-grading results. To answer RQ1, we extracted students’ last written prompt and graded them manually to assess AI auto-grading. After receiving 483 unique student-written prompts, two researchers used the same rubric to grade the students’ written prompts (Table 1). They first randomly selected 15% of the prompt data and conducted human grading based on the rubric. After that, the two researchers met to address the conflicts and iteratively refined the detailed grading book. Then they randomly selected another 15% of the prompt data and both two researchers graded it again; the inter-rater reliability for this round reached higher than 0.92 (almost perfect agreement following (mchugh2012interrater)). Then the two researchers met again to resolve the conflicts and graded the rest. We used human grading results for RQ2. For RQ3, we conducted a thematic analysis on student answers. One hundred and thirty-one students finished answering the self-reflection questions. Two researchers conducted a coding book iteratively to analyze their answers and reported the main themes.
@@ 4.3 Data Collection and Analysis
Table 2: Grading Accuracy of the Auto-Grader
4.4 RQ1: What is the accuracy of AI-based grading for students’ written prompts?
@@
The auto-grader achieved a high accuracy (higher than 0.9) in Relevance, Background, and Elaboration dimensions. For the inaccurate cases in these dimensions, one pattern is that the auto-grader tended to weigh heavily on some keywords but ignored other keywords (nguyen2023evaluating). For instance, one prompt was “You are preparing for a geography exam and have just learned the basic concepts of the water cycle Answer in Sanskrit.” The auto-grader gave it a False in Relevance and explained as “The core of the problem lies in language choice and has nothing to do with the specific content of the water cycle.”.
@@
The accuracy for Conciseness and No Direct Answer was 0.93 and 0.88. The conflicted cases for Conciseness mostly included two types. First, the auto-grader focused on spelling errors and therefore graded them as False, while human graders were more tolerant about this. Second, prompts containing background information might be seen as non-concise by the auto-grader. For example, one student-written prompt was “Can you tell me in detail about your extracurricular knowledge about cells besides the definition of cells?”. The auto-grader graded Conciseness as False and explained it as “The question is relatively long and the sentences are not concise enough. It could be more refined.”
@@
As for No Direct Answer, sometimes the student asked a general question about “How to solve a two-variable linear equation?”, but the auto-grader treated this as the student tried to ask for a direct answer indirectly and labeled it as False. On the other hand, sometimes the student asked for “list the steps to solve this equation 10x+4y=3,-2x+10y=4”, but auto-grader thought this prompt “did not directly ask about the values of x and y, but rather asks for a solution.” and gave it True in No Direct Answer instead.
@@
Auto-grader received the lowest accuracy in Purpose dimension (0.85). The main reason for false positive cases in this dimension is that the auto-grader might over-generate the keywords in a student-written prompt, and consider that as providing a clear purpose. For instance, it counted “I am a junior high school student. Please help me solve this problem using junior high school knowledge.” as True in Purpose dimension and explained it as “the student’s goal is to solve this problem.” However, simply asking for “solving this problem” without specifying the type or domain does not align with our rubric (“specific and clear purpose, e.g. explain the math concepts involved”).
@@
For true negative cases in Purpose, sometimes the auto-grader mixed the criteria of Purpose and No Answer, and therefore considered those involving asking for an answer as False. Another reason is that auto-grader sometimes requires too detailed purpose, which is outside of our targeted scope. For example, one explanation to an AI-graded False case (“I want to know about DNA”) is “Although DNA is mentioned in this question, the purpose is not clear. Students should be more specific about what they want to learn about DNA.” However, as the scenario is novice learners who only finished the first biology class about cells, the auto-grader should not expect the prompt purpose to be that detailed.
4.5 RQ2: How does this activity influence students’ prompting ability and confidence in using AI for future learning?
@@
Students answered four 5-scale Likert questions as a pre-survey before doing the pre-test. After finishing the learning activities, they answered the last question in the pre-survey again. Specific questions and the student response distribution are shown in Figure 3.
@@
(Applied human-graded Q2 0.93 0.65 0.94 0.12 0.59 - -
scores, Yes-1/No-0) Q3 0.89 0.75 0.95 0.11 0.52 0.64 -
@@
Ninety-eight students completed both the pre- and post-test. However, no significant increment (p = .377) was found in the post-test (Mean = 4.4, Median = 5.0, SD = 0.96, full score = 6.0) due to the ceiling effect in the pre-test (Mean = 4.4, Median = 5.0, SD = 1.04). In other words, students had the basic skills to identify the prompt quality from a conceptual standpoint, although their prompt writing skills remain varied.
4.6 RQ3: How do students perceive the benefits and challenges of this experience?